OpenStudy (kj4uts):

Determine the average rate of change of the function f(x)=-2x^2+5 over the interval from x=1 to x=4.

OpenStudy (kj4uts):

Do I plug this in like this: -2(1)^2+5=3 -2(4)^2+5=-27

OpenStudy (kj4uts):

My choices are: -10 10 -8 8 None of the others

Nnesha (nnesha):

average rate of change = slope

Nnesha (nnesha):

you pluggd 1 and 4 to find y values now use slope formula to find average rate of change \[\huge\rm \frac{ y_2 -y_1 }{ x_2 -x_1 }\]

OpenStudy (kj4uts):

-10

Nnesha (nnesha):

:=) looks good